OpenStudy (anonymous):

Ingrid is making a quilt using squares that measure 5 in. on a side. What is the length of a diagonal of one of the quilt squares? Round to the nearest tenth

OpenStudy (anonymous):

use Pythagorean Theorem since it's a squares... a^2 + b^2 = c2 .....5^2 + 5^2 = c^2 25 + 25 = c^2 50 = c^2 c = 7.0710... c = 7.1in

OpenStudy (anonymous):

oh ok thanks :)

OpenStudy (anonymous):

You're welcome! : )